OpenStudy (anonymous):

Written equation I need help with. "Saidhari bought a toaster oven for $75. This price was 1/4 off the list price" What was the list price?

OpenStudy (anonymous):

I see you came back.

OpenStudy (butterflydreamer):

okay so, Let list price = x We know the toaster costs $75 which is 1/4 OFF the list price... This means: 75 = x - 1/4x 75 = 3/4x So then divide both sides by 3/4 to find "x" which is your list price. x = ?

OpenStudy (anonymous):

you gotta keep a toaster handy to warm your bread, you feel me

OpenStudy (anonymous):

$1.00? Because I got 100.

OpenStudy (butterflydreamer):

remember, 75 as in $75.. which is 3/4 of the original amount. SO, original amount is bigger than $75 so... list price = $100 not $1.. or else the toaster would be 75 cents (if you took off 1/4) LOL

OpenStudy (anonymous):

Thank you, that was the correct answer.

OpenStudy (butterflydreamer):

Look, we can always double check by working backwards. So list price is $100 1/4 of the list price is 100 * 1/4 = $25 So toaster after 1/4 of list price is taken off is : List price - 1/4 list price Toaster after 1/4 of list price is taken off is : $100 - $25 = $75 which is correct :)! You're welcome!